# Gaussian Head Avatar: Ultra High-fidelity Head Avatar via Dynamic Gaussians

### v2

@@ 4.1 Avatar Representation
Next, we explain the process of adding expression-related changes to the neutral Gaussian model through the dynamic generator $\boldsymbol{\Phi}$ as described in Eqn. 2 in detail.
@@
Positions $X^{\prime}$ of the Gaussians. Expressions bring about the geometric deformation of the neutral model, which is modeled as the displacements of the Gaussian points. Specifically, we predict the displacements respectively controlled by the expression and the head pose in the canonical space through two different MLPs: $\boldsymbol{f}_{def}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{def}^{pose}\in\boldsymbol{\Phi}$. Then, we add them to the neutral positions.
@@
$\begin{split}X^{\prime}=\boldsymbol{X}_{0}+\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{def}^{exp}(\boldsymbol{X}_{0},\theta)\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{def}^{pose}(\boldsymbol{X}_{0},\beta).\end{split}$ (3)
@@
$\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ represent the extent to which the points are affected by the expression or the head pose respectively. Without decoupling but just using the expression coefficients as the global condition [16] which also controls the shoulders and upper body, will produce jittering results during the animation. Here, we assume that the Gaussian points closer to 3D landmarks are more affected by the expression coefficients and less affected by the head pose, while the opposite is true for the Gaussian points far away. Specifically, The 3D landmarks $\boldsymbol{P}_{0}$ of the canonical model are first estimated through the 3DMM model in the data preprocessing and then optimized in the initialization stage 4.3. Then for each Gaussian point, we calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ as follows:
@@
$\begin{split}\lambda_{exp}(x)=\left\{\begin{array}[]{ll}1,&dist(x,\boldsymbol{P}_{0})<t_{1}\\ \frac{t_{2}-dist(x,\boldsymbol{P}_{0})}{t_{2}-t_{1}},&dist(x,\boldsymbol{P}_{0})\in[t_{1},t_{2}]\\ 0,&dist(x,\boldsymbol{P}_{0})>t_{2}\end{array}\right.\end{split}$
@@
with $\lambda_{pose}(x)=1-\lambda_{exp}(x)$. And $x\in\mathbf{X}_{0}$ denotes the position of one neutral Gaussian. $dist(x,\boldsymbol{P}_{0})$ denotes the minimum distance from the point $x$ to the 3D landmarks $\boldsymbol{P}_{0}$. $t_{1}=0.15$ and $t_{2}=0.25$ are predefined hyperparameters when the length of the head is set to approximately $1$.
@@
Color $C^{\prime}$ of the Gaussians. Modeling the dynamic details typically requires dynamic color that changes with expressions. As we do not pre-define the neutral value in Eqn. 2, the color are directly predict by two color MLPs: $\boldsymbol{f}_{col}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{col}^{pose}\in\boldsymbol{\Phi}$:
@@
$\begin{split}C^{\prime}=\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{col}^{exp}(\boldsymbol{F}_{0},\theta)\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{col}^{pose}(\boldsymbol{F}_{0},\beta).\end{split}$ (4)
@@
Rotation, Scale and Opacity $\{Q^{\prime},S^{\prime},A^{\prime}\}$ of the Gaussians. These three attributes also dynamic, thereby modeling some detailed expressions-related appearance changes. Here, we just use another two attribute MLPs $\boldsymbol{f}_{att}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{att}^{pose}\in\boldsymbol{\Phi}$ to predict their shift from the neutral value.
@@
$\begin{split}\{Q^{\prime},S^{\prime},A^{\prime}\}=\{\mathbf{Q}_{0},\mathbf{S}_{0},\mathbf{A}_{0}\}\\ +\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{att}^{exp}(\boldsymbol{F}_{0},\theta)\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{att}^{pose}(\boldsymbol{F}_{0},\beta).\end{split}$ (5)
@@
Finally, we apply rigid rotations and translations $T(\cdot)$ to the Gaussians, transforming them from the canonical space to the world space. Note, the transformation is only implemented for directional variables: $\{X^{\prime},Q^{\prime}\}$, while the multi-channel color, the scale and the opacity $\{C^{\prime},S^{\prime},A^{\prime}\}$ are not directional thus remain unchanged.
$\{X,Q\}=T(\{X^{\prime},Q^{\prime}\},\beta),$ (6)
$\{C,S,A\}=\{C^{\prime},S^{\prime},A^{\prime}\}.$ (7)
@@ 4.2 Training
In this part, we explain the training pipeline of the Gaussian head avatar 4.1 and the loss function. In each iteration, we first generate the expression conditioned 3D Gaussians as Eqn. 2. Then we render a 32-channel image with 512 resolution $I_{C}\in\mathbb{R}^{512\times 512\times 32}$ referring to Eqn. 1. After that we feed the image to a super resolution network $\boldsymbol{\Psi}$ to generate a 2048 resolution RGB image $I_{hr}\in\mathbb{R}^{2048\times 2048\times 3}$, such that more details are recovered and noise caused by uneven ambient light or camera chromatic aberration in the training data will be filtered out [64, 49].
During training, we jointly optimize all the learnable parameters mentioned above in bold, including the neutral Gaussians: $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, the dynamic generator: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose},\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\}$, and the super resolution network $\boldsymbol{\Psi}$. For the loss function, we only use the foreground RGB images $I_{gt}$ as supervision to construct an L1 loss and a VGG perceptual loss [70] between the generated images $I_{hr}$ and the ground truth $I_{gt}$. Besides, we encourage the first three channels of the 32-channel feature image $I_{C}$ to be RGB channels, which is ensured by a L1 loss term. The total loss is:
$\begin{split}\mathcal{L}=||I_{hr}-I_{gt}||_{1}+\lambda_{vgg}VGG(I_{hr},I_{gt})\\ +\lambda_{lr}||I_{lr}-I_{gt}||_{1},\end{split}$ (8)
@@ 4.3 Geometry-guided Initialization
Mesh Guidance Model. Specifically, we first construct a MLP $\boldsymbol{f}_{sdf}$ to represent a signed distance field. In addition, this network will also output the corresponding feature vector of each point, which is used for predicting the point color. It can be formulated as:
${s,\eta}=\boldsymbol{f}_{sdf}(x),$ (9)
with $s$ denotes the SDF value, $\eta$ denotes the feature vector and $x$ denotes the point position. Then through (DMTet) [51], we can differentially extract a mesh with vertices $X$, per-vertex feature vectors $F$ and its faces. We also predict the per-vertex 32-channel color as Eqn. 4 by the two color MLPs $\boldsymbol{f}_{col}^{exp}$ and $\boldsymbol{f}_{col}^{pose}$. In parallel, we construct the two deformation MLPs: $\boldsymbol{f}_{def}^{exp}$ and $\boldsymbol{f}_{def}^{pose}$ as described in Sec. 4.1 to predict the displacements and add them to the vertex positions. This process is similar to Eqn 3 above, with the Gaussian positions $\boldsymbol{X}_{0}$ replaced by the vertex positions $X$. Finally we also apply rigid rotations and translations to the deformed mesh, transforming it to the world space and render the deformed mesh into an image $I$ and a mask $M$ through differentiable rasterization [44] according to the camera parameters $\mu$.
Loss Function and Training. Next, we can construct the RGB loss and the silhouette loss to train the guidance model:
@@
We also use the estimated 3D facial landmarks $P_{gt}$ as described in Sec. 3 to provide rough guidance for the expression deformation MLP. Specifically, we input the neutral 3D landmarks $\boldsymbol{P}_{0}$ into the expression deformation MLP to predict the expression conditioned landmarks $P$:
@@
$\mathcal{L}_{def}=||P-P_{gt}||_{2},$ (13)
Besides, we introduce three constraints: (1) a regular term $\mathcal{L}_{offset}$ to punish all non-zero displacements to prevent the two deformation MLPs from learning a global constant offset [63], (2) a regular term $\mathcal{L}_{lmk}$ to limit the SDF value at the 3D landmarks to be close to zero, such that the landmarks are located on the surface of the mesh, (3) a Laplacian term $\mathcal{L}_{lap}$ for maintaining the extracted mesh smooth to a certain extent. Overall, the total loss function is formulated as:
$\begin{split}\mathcal{L}=\mathcal{L}_{RGB}+\lambda_{sil}\mathcal{L}_{sil}+\lambda_{def}\mathcal{L}_{def}+\\ \lambda_{offset}\mathcal{L}_{offset}+\lambda_{lmk}\mathcal{L}_{lmk}+\lambda_{lap}\mathcal{L}_{lap},\end{split}$ (14)
with $\lambda$ denoting the weights of each term, which are set as follows: $\lambda_{sil}=0.1$, $\lambda_{def}=1$, $\lambda_{offset}=0.01$, $\lambda_{lmk}=0.1$ and $\lambda_{lap}=100$. We jointly optimize the MLPs mentioned above with the neutral 3D landmarks $\boldsymbol{P}_{0}$ jointly until all MLPs are converged.
Parameters Transfer. Finally, we use the roughly trained mesh guidance model to initialize the Gaussian model. Specifically, we extract the neutral mesh with vertices $X$ and per-vertex features $F$ through DMTet, and directly assign their values to the neutral positions $\boldsymbol{X}_{0}=X$ and the per-vertex feature vectors $\boldsymbol{F}_{0}=F$ of the neutral Gaussians respectively. Then, we retain all the four optimized MLPs: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose}\}$ for the Gaussian model. For the other neutral attributes: rotation, scale and opacity, we adopt the original initialization strategy in Gaussian Splatting [26]. And the parameters of the two attribute MLPs: $\{\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\}$ and the super resolution network $\boldsymbol{\Psi}$ are just randomly initialized.

### v3

@@ 3.1.1 Head Avatar
Next, we explain the process of adding expression-related changes to the neutral Gaussian model through the dynamic generator $\boldsymbol{\Phi}$ as described in Eqn. 2 in detail. We use superscript (’) to denote the parameters of Gaussians which are changed by MLPs in the canonical space but not transformed to the world space by head pose $\beta_{t}$.
@@
Positions $X^{\prime}_{t}$ of the Gaussians at time step $t$. Expressions bring about the geometric deformation of the neutral model, which is modeled as the displacements of the Gaussian points. Specifically, we predict the displacements respectively controlled by the expression and the head pose in the canonical space through two different MLPs: $\boldsymbol{f}_{def}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{def}^{pose}\in\boldsymbol{\Phi}$. Then, we add them to the neutral positions.
@@
$\begin{split}X^{\prime}_{t}=\boldsymbol{X}_{0}+\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{def}^{exp}(\boldsymbol{X}_{0},\theta_{t})\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{def}^{pose}(\boldsymbol{X}_{0},\beta_{t}).\end{split}$ (3)
@@
$\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ represent the extent to which the points are affected by the expression or the head pose respectively. Without decoupling but just using the expression coefficients as the global condition [68] which also controls the shoulders and upper body, will produce jittering results during the animation. Here, we assume that the Gaussian points closer to 3D landmarks are more affected by the expression coefficients and less affected by the head pose, while the opposite is true for the Gaussian points far away. Specifically, The 3D landmarks $\boldsymbol{P}_{0}$ of the canonical model are first estimated through the 3DMM model in the data preprocessing and then optimized in the initialization stage 3.3. Then for each Gaussian point, we calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ as follows:
@@
$\begin{split}\lambda_{exp}(x)=\left\{\begin{array}[]{ll}1,&dist(x,\boldsymbol{P}_{0})<t_{1}\\ \frac{t_{2}-dist(x,\boldsymbol{P}_{0})}{t_{2}-t_{1}},&dist(x,\boldsymbol{P}_{0})\in[t_{1},t_{2}]\\ 0,&dist(x,\boldsymbol{P}_{0})>t_{2},\end{array}\right.\end{split}$
@@
with $\lambda_{pose}(x)=1-\lambda_{exp}(x)$. And $x\in\mathbf{X}_{0}$ denotes the position of one neutral Gaussian. $dist(x,\boldsymbol{P}_{0})$ denotes the minimum distance from the point $x$ to the 3D landmarks $\boldsymbol{P}_{0}$. $t_{1}=0.15$ and $t_{2}=0.25$ are predefined hyperparameters when the length of the head is set to approximately $1$. For the hairs, we set $\lambda_{exp}(x)=0$ since the hairs are hardly affected by the expression coefficients.
@@
Color $C^{\prime}_{t}$ of the Gaussians at time step $t$. Modeling the dynamic details typically requires dynamic color that changes with expressions. As we do not pre-define the neutral value in Eqn. 2, the color are directly predict by two color MLPs: $\boldsymbol{f}_{col}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{col}^{pose}\in\boldsymbol{\Phi}$:
@@
$\begin{split}C^{\prime}_{t}=\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{col}^{exp}(\boldsymbol{F}_{0},\theta_{t})\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{col}^{pose}(\boldsymbol{F}_{0},\beta_{t}).\end{split}$ (4)
@@
Rotation, Scale and Opacity $\{Q^{\prime}_{t},S^{\prime}_{t},A^{\prime}_{t}\}$ of the Gaussians at time step $t$. These three attributes also dynamic, thereby modeling some detailed expressions-related appearance changes. Here, we just use another two attribute MLPs $\boldsymbol{f}_{att}^{exp}\in\boldsymbol{\Phi}$ and $\boldsymbol{f}_{att}^{pose}\in\boldsymbol{\Phi}$ to predict their shift from the neutral value.
@@
$\begin{split}\{Q^{\prime}_{t},S^{\prime}_{t},A^{\prime}_{t}\}=\{\mathbf{Q}_{0},\mathbf{S}_{0},\mathbf{A}_{0}\}\\ +\lambda_{exp}(\boldsymbol{X}_{0})\boldsymbol{f}_{att}^{exp}(\boldsymbol{F}_{0},\theta_{t})\\ +\lambda_{pose}(\boldsymbol{X}_{0})\boldsymbol{f}_{att}^{pose}(\boldsymbol{F}_{0},\beta_{t}).\end{split}$ (5)
3.1.2 Dynamic Hair
Next, we explain the process of the changes in hair with head motion in detail. We assume the head and the hair maintain still at time step 00, which means that hairs do not have extra deformation due to the inertia, the hair positions $X_{0}=X_{-1}=X_{-2}$, and the head poses $\beta_{0}=\beta_{-1}=\beta_{-2}$.
Hair deformation. The deformation of hair is related to the previous hair position $X_{t-1}$ which can get from $X^{\prime}_{t-1}$ and $\beta_{t-1}$, hair speed which can get from $X^{\prime}_{t-1}$, $X^{\prime}_{t-2}$, $\beta_{t-1}$, and $\beta_{t-2}$, head pose $\beta_{t}$, and head movement speed which can get from $\beta_{t}$ and $\beta_{t-1}$. We use a MLP $\boldsymbol{f}_{def}^{hair}\in\boldsymbol{\Phi}$ to predict the extra hair deformation.
[…]
Here, $X^{\prime}_{t-1,t-2}$ and $\beta_{t,t-1,t-2}$ represent respectively the $X^{\prime}$ at time step ${t-1,t-2}$ and the $\beta$ at time step ${t,t-1,t-2}$. $\lambda_{hair}(\cdot)$ represents the extent to which the points of the hair are affected by the head pose. The strength of the hair deformation is related to the distance between the hairs and the scalp. We assume that the Gaussian points closer to the scalp are more affected by the head pose, while the opposite is true for the Gaussian points far away. Specifically, The scalp $\boldsymbol{P}_{1}$ of the canonical model are first estimated through the FLAME model [24] in the data preprocessing. Then for each Gaussian point, we calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ as follows:
$\begin{split}\lambda_{hair}(x)=\left\{\begin{array}[]{ll}1,&dist(x,\boldsymbol{P}_{1})<t_{3}\\ \frac{t_{4}-dist(x,\boldsymbol{P}_{1})}{t_{4}-t_{3}},&dist(x,\boldsymbol{P}_{1})\in[t_{3},t_{4}]\\ 0,&dist(x,\boldsymbol{P}_{1})>t_{4},\end{array}\right.\end{split}$
with $x\in\mathbf{X}_{0}$ denotes the position of one neutral Gaussian. $dist(x,\boldsymbol{P}_{1})$ denotes the minimum distance from the point $x$ to the scalp $\boldsymbol{P}_{1}$. $t_{3}=0.05$ and $t_{4}=0.15$ are predefined hyperparameters when the length of the head is set to approximately $1$.
Hair attribution. The Gaussian properties of hair should also undergo corresponding changes after deformation. Here we use two MLPs $\boldsymbol{f}_{col}^{hair}$ and $\boldsymbol{f}_{att}^{hair}$ $\in\boldsymbol{\Phi}$ to predict the extra changes in the Gaussian properties of hair.
$\begin{split}C^{\prime}_{t}=C^{\prime}_{t}+\boldsymbol{f}_{col}^{hair}(\boldsymbol{X}_{0},X^{\prime}_{t-1,t-2},\beta_{t,t-1,t-2}),\end{split}$ (7)
[…]
@@ 3.1.2 Dynamic Hair
Finally, we apply rigid rotations and translations $T(\cdot)$ to the Gaussians, transforming them from the canonical space to the world space. Note, the transformation is only implemented for directional variables: $\{X^{\prime}_{t},Q^{\prime}_{t}\}$, while the multi-channel color, the scale and the opacity $\{C^{\prime}_{t},S^{\prime}_{t},A^{\prime}_{t}\}$ are not directional thus remain unchanged.
$\{X_{t},Q_{t}\}=T(\{X^{\prime}_{t},Q^{\prime}_{t}\},\beta_{t}),$ (9)
$\{C_{t},S_{t},A_{t}\}=\{C^{\prime}_{t},S^{\prime}_{t},A^{\prime}_{t}\}.$ (10)
@@ 3.2 Training
In this part, we explain the training pipeline of the HHAvatar 3.1 and the loss function. In each iteration, we first generate the expression conditioned 3D Gaussians as Eqn. 2. Then, given a camera view, we render a 32-channel image with 512 resolution $I_{C}\in\mathbb{R}^{512\times 512\times 32}$ referring to Eqn. 1. After that we feed the image to a super resolution network $\boldsymbol{\Psi}$ to generate a 2048 resolution RGB image $I_{hr}\in\mathbb{R}^{2048\times 2048\times 3}$, such that more details are recovered and noise caused by uneven ambient light or camera chromatic aberration in the training data will be filtered out [72, 79]. Meanwhile, the expressive Gaussians are also rendered to a segmentation map for separating the hair from the head.
During training, we jointly optimize all the learnable parameters mentioned above in bold, including the neutral Gaussians: $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, the dynamic generator: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose},\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\boldsymbol{f}_{def}^{hair},\boldsymbol{f}_{col}^{hair},\boldsymbol{f}_{att}^{hair}\}$, and the super resolution network $\boldsymbol{\Psi}$. Our training loss consists of two parts: the reconstruction loss $\mathcal{L}_{recon}$ and the mask loss $\mathcal{L}_{mask}$:
$\begin{split}\mathcal{L}=\mathcal{L}_{recon}+\lambda_{mask}\mathcal{L}_{mask}.\end{split}$ (11)
Reconstruction loss. For the reconstruction loss function, we only use the foreground RGB images $I_{gt}$ as supervision to construct an L1 loss and a VGG perceptual loss [96] between the generated images $I_{hr}$ and the ground truth $I_{gt}$. Besides, we encourage the first three channels of the 32-channel feature image $I_{C}$ to be RGB channels, which is ensured by a L1 loss term. The total loss is:
$\begin{split}\mathcal{L}_{recon}=||I_{hr}-I_{gt}||_{1}+\lambda_{vgg}VGG(I_{hr},I_{gt})\\ +\lambda_{lr}(||I_{lr}-I_{gt}||_{1}+\lambda_{vgg}VGG(I_{lr},I_{gt})),\end{split}$ (12)
@@
Mask loss. To separate the points of the head and the hair, we introduce the occlusion perception module to use the head mask and the hair mask to supervise the head part and the hair part. We first use face-parsing [95] to obtain the mask of the hair and the head. Then, we set the Gaussian color of the head and the hair to $(1.0,0.0,0.0)$ and $(0.0,1.0,0.0)$ respectively. Finally, we can get the predicted color by rendering to obtain the head mask and the hair mask, which can avoid supervision of the parts covered by hair.
$\begin{split}\mathcal{L}_{mask}=||M^{\prime}_{head}-M_{head}||_{1}+||M^{\prime}_{hair}-M_{hair}||_{1},\end{split}$ (13)
with $M^{\prime}_{head}$ and $M^{\prime}_{hair}$ denoting the predicted head mask and the predicted hair mask respectively, $M_{head}$ and $M_{hair}$ denoting the head mask and the hair mask from face-parsing [95]. We set the weight $\lambda_{mask}=0.1$.
@@ 3.3 Geometry-guided Initialization
Mesh Guidance Model. Specifically, we first construct two MLPs $\boldsymbol{f}_{sdf\_head}$ and $\boldsymbol{f}_{sdf\_hair}$ to represent two signed distance fields for the head and the hair respectively. In addition, this network will also output the corresponding feature vector of each point, which is used for predicting the point color. It can be formulated as:
${s_{head},\eta_{head}}=\boldsymbol{f}_{sdf\_head}(x),$ (14)
${s_{hair},\eta_{hair}}=\boldsymbol{f}_{sdf\_{hair}}(x),$ (15)
with $s_{head}$ and $s_{hair}$ denote the SDF value of the head and the hair respectively, $\eta_{head}$ and $\eta_{hair}$ denote the feature vector of the head and the hair respectively and $x$ denotes the point position. Then through DMTet [32], we can differentially extract the mesh with vertices $X$, per-vertex feature vectors $F$, and its faces for the head and the hair respectively, and we merge the mesh of the head and the hair to obtain the complete mesh of the whole head. We also predict the per-vertex 32-channel color as Eqn. 4 by the two color MLPs $\boldsymbol{f}_{col}^{exp}$ and $\boldsymbol{f}_{col}^{pose}$ for the head and the hair. In parallel, we construct the two deformation MLPs: $\boldsymbol{f}_{def}^{exp}$ and $\boldsymbol{f}_{def}^{pose}$ as described in Sec. 3.1 to predict the displacements and add them to the vertex positions. This process is similar to Eqn 3 above, with the Gaussian positions $\boldsymbol{X}_{0}$ replaced by the vertex positions $X$. Finally, we also apply rigid rotations and translations to the deformed mesh, transforming it to the world space and rendering the deformed mesh into an image $I$, a mask $M$, a head mask $M^{\prime}_{head}$, and a hair mask $M^{\prime}_{hair}$ through differentiable rasterization [97] according to the camera parameters $\mu$. Note that during the geometry-guided initialization, we temporarily ignore the real physical motion of the hair (i.e., we do not consider the additional motion of hair due to inertia).
Loss Function and Training. Next, we can first construct the RGB loss and the silhouette loss to train the guidance model:
@@
We also use the estimated 3D facial landmarks $P_{gt}$ to provide rough guidance for the expression deformation MLP. Specifically, we input the neutral 3D landmarks $\boldsymbol{P}_{0}$ into the expression deformation MLP to predict the expression conditioned landmarks $P$:
@@
$\mathcal{L}_{def}=||P-P_{gt}||_{2}.$ (19)
Besides, we introduce three constraints: (1) a regularization term $\mathcal{L}_{offset}$ to punish all non-zero displacements to prevent the two deformation MLPs from learning a global constant offset [71], (2) a regularization term $\mathcal{L}_{lmk}$ to limit the SDF value at the 3D landmarks to be close to zero, such that the landmarks are located on the surface of the mesh, (3) a Laplacian term $\mathcal{L}_{lap}$ for maintaining the extracted mesh smooth to a certain extent. To obtain the signed distance field of the head and the hair respectively, we also use the head mask and the hair mask to supervise the head part and the hair part:
$\begin{split}\mathcal{L}_{mask}=IOU((1-M_{o})\cdot M^{\prime}_{head},M_{head})\\ +IOU(M_{o}\cdot M^{\prime}_{hair},M_{hair}),\end{split}$ (20)
$M_{o}=D_{hair}<D_{head},$ (21)
with $M^{\prime}_{head}$ and $M^{\prime}_{hair}$ denoting the predicted head mask and the predicted hair mask respectively, $M_{head}$ and $M_{hair}$ denoting the head mask and the hair mask from face-parsing [95] respectively, and $D_{head}$ and $D_{hair}$ denoting the predicted head depth and the predicted hair depth respectively.
Overall, the total loss function in the initialization stage is formulated as:
$\begin{split}\mathcal{L}=\mathcal{L}_{RGB}+\lambda_{sil}\mathcal{L}_{sil}+\lambda_{def}\mathcal{L}_{def}\\ +\lambda_{offset}\mathcal{L}_{offset}+\lambda_{lmk}\mathcal{L}_{lmk}\\ +\lambda_{lap}\mathcal{L}_{lap}+\lambda_{mask}\mathcal{L}_{mask},\end{split}$ (22)
with $\lambda$ denoting the weights of each term, which are set as follows: $\lambda_{sil}=0.1$, $\lambda_{def}=1$, $\lambda_{offset}=0.01$, $\lambda_{lmk}=0.1$, $\lambda_{lap}=100$, and $\lambda_{mask}=0.1$. We jointly optimize the MLPs mentioned above with the neutral 3D landmarks $\boldsymbol{P}_{0}$ jointly until all MLPs are converged.
Parameters Transfer. Finally, we use the roughly trained mesh guidance model to initialize the Gaussian model. Specifically, we extract the neutral mesh with vertices $X$ and per-vertex features $F$ through DMTet [32] for the head and the hair respectively, and directly assign their values to the neutral positions $\boldsymbol{X}_{0}=X$ and the per-vertex feature vectors $\boldsymbol{F}_{0}=F$ of the neutral Gaussians respectively. Then, we retain all the four optimized MLPs: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose}\}$ for the Gaussian model. For the other neutral attributes: rotation, scale and opacity, we adopt the original initialization strategy in Gaussian Splatting [9]. And the parameters of the two attribute MLPs: $\{\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose},\boldsymbol{f}_{def}^{hair},\boldsymbol{f}_{col}^{hair},\boldsymbol{f}_{att}^{hair}\}$ and the super resolution network $\boldsymbol{\Psi}$ are just randomly initialized.